8. Consider a newsvendor environment where demand is expected to be uniformly distributed between 1000 and 4000 units. The variable production cost is $20 per unit and the sales price is $50 per unit. In addition, the producer must pay $5 to discard any unsold units. How many units should be produced? SHOW WORK

a. 3000

b. 2636

c. 2500

d. 3182

e. 3308
Answer #1

Underage cost, Cu = Sales price - Variable production cost = 50 - 20 = $ 30

Overage cost, Co = Variable production cost + Cost to discard unsold units = 20 + 5 = $ 25

Optimal service level, s = Cu/(Cu+Co) = 30/(30+25) = 0.5455

Number of units to be produced = a + s*(b-a) = 1000 + 0.5455*(4000-1000) = 2636

ANSWER: b. 2636